OpenStudy (anonymous):

9^8x = 27 How do you do this? I converted it to log base 9 of 27 = 8x but that didn't really help me.

OpenStudy (phi):

I notice that 9 is 3^2 and 27 is 3^3 so replace 9 with 3^2 and 27 with 3^3

OpenStudy (phi):

\[ (3^2)^{8x} = 3^3 \]

OpenStudy (anonymous):

What do you do after that? Sorry I really don't get logs.

OpenStudy (phi):

I would use the rule \[ (x^a)^b= x^{ab} \] to rewrite the left side then notice that you have the same bases on both sides. I would think the exponents must equal each other for the equation to be true. or take the \(\log_3 \) of both sides

OpenStudy (anonymous):

Thanks, that helped a lot for all of my log problems!

OpenStudy (phi):

for this problem you should get 16x =3 and x= 3/16

OpenStudy (precal):

this is another way to do it